OpenStudy (anonymous):

what would be the equation after completing the square and factoring? x^2+5x+8=4

OpenStudy (x3_drummerchick):

first set the equation equal to 0 by subtracting 4 \[x ^{2} + 5x + 4\]

OpenStudy (anonymous):

ok whats next?

OpenStudy (x3_drummerchick):

factoring

OpenStudy (anonymous):

it would be \[x=-4,-1\]

OpenStudy (x3_drummerchick):

(x + 4 ) ( x + 1 )

OpenStudy (x3_drummerchick):

correct

OpenStudy (x3_drummerchick):

because you're setting both (x+4)= 0 and (x+1)=0

OpenStudy (anonymous):

yea but how would you write the equation

OpenStudy (anonymous):

because first you need to complete the square and then factor and when that is done you have to write it in an equation

OpenStudy (x3_drummerchick):

was the question said to write the equation, then complete the square and factoring?

OpenStudy (anonymous):

no it said what would be the equationafter competing the square and factoring \[x^2+5x+8=4\]

zepdrix (zepdrix):

So when drumchica got to (x+4)(x+1)=0, that's where you would stop. That is the factored form they're looking for. Are you having any trouble understanding the steps up to that point?